I got to find the slop of the line through the following pairs of points


(-5 , -3 ) and (-5, 2)

Answer by nik_consult(24)

(-5 , -3 ) and (-5, 2)
slope, m= y2-y1 / x2-x1 ; y2=2, y1=-3, x2=-5 and x1=-5
= 2 - (-3) / -5 - (-5)
= 2 + 3 / 0
= 5/ 0
= Undefined